# Neutron Star-White Dwarf Merger as One Possible Optional Source of Kilonova-like Emission: Implications for GRB 211211A

### v1

@@ Abstract
Long-duration GRB 211211A, which lacked an associated supernova at such a low redshift $z=0.076$, but was associated with a possible kilonova emission, has attracted great attention. The neutron star-white dwarf (NS-WD) merger is proposed as a possible progenitor of GRB 211211A, and it could naturally explain the long duration of the prompt emission. However, the NS-WD merger is not an ideal site for producing heavy elements via r-process nucleosynthesis. In this Letter, we investigate the heavy elements produced in NS-WD mergers based on numerical simulations of nucleosynthesis via SkyNet, and then calculate the resulting kilonova-like emission to compare with the solidly observed case of possible kilonova emission associated with GRB 211211A. By adopting three models (i.e., Model-A, Model-B, and Model-C) from Kaltenborn et al. (2023) at different temperatures ($T=4$ GK, 5 GK, and 6 GK), which are treated as free parameters, we find that the mass number of the heaviest element produced in our simulations is less than 90 ($A<90$). Moreover, by comparing the calculated kilonova-like emission with the afterglow-subtracted observations of the possible kilonova associated with GRB 211211A, it is found that the merger of NS and WD cannot be ruled out as the origin of GRB 211211A to produce the possible kilonova emission if the remnant of the WD-NS merger is a supramassive or stable magnetar. Otherwise, it is difficult to explain the early possible kilonova emission following GRB 211211A by invoking the merger of a WD and an NS.
Gamma-ray bursts, Nucleosynthesis, Neutron stars, White dwarf stars
@@ 1 Introduction
Gamma-ray bursts (GRBs), the most luminous electromagnetic explosions in the universe, are traditionally classified into two categories based on their duration: short-duration GRBs (SGRBs; $T_{90}<2s$) and long-duration GRBs (LGRBs; $T_{90}>2s$) (Kouveliotou et al., 1993). Robust associations of the underlying supernovae (SNe) with some LGRBs suggest that they are likely related to the collapse of massive stars (Woosley, 1993; Galama et al., 1998; Stanek et al., 2003; Pian et al., 2006; Fruchter et al., 2006; Woosley & Bloom, 2006). In contrast, some SGRBs associated with nearby host galaxies with little star formation, and without an associated SN (Tanvir et al., 2005; Fox et al., 2005; Hjorth et al., 2005; Berger et al., 2005), suggesting that the SGRBs are possibly originated from the merger of binary compact stars, such as binary neutron star (BNS) mergers (Paczynski, 1986; Eichler et al., 1989) and neutron star-black hole (NS-BH) mergers (Paczynski, 1991). On the other hand, such binary systems are also considered to be potential sources of gravitational-wave radiation and kilonova emission (Berger, 2014; Zhang, 2018). The first “smoking gun” evidence to confirm the BNS merger as the source of SGRBs, gravitational-wave radiation, and kilonova emerged from the direct detection of GW 170817 via Advanced LIGO and Virgo, which was associated with electromagnetic counterparts, such as GRB 170817A and AT 2017gfo (Abbott et al., 2017; Alexander et al., 2017; Arcavi et al., 2017; Chornock et al., 2017; Coulter et al., 2017; Covino et al., 2017; Drout et al., 2017; Evans et al., 2017; Goldstein et al., 2017; Kasen et al., 2017; Kasliwal et al., 2017; Pian et al., 2017; Savchenko et al., 2017; Shappee et al., 2017; Smartt et al., 2017; Tanaka et al., 2017; Tanvir et al., 2017; Zhang et al., 2018).
@@
Excepting the confirmed event GW 170817/ GRB 170817A/ AT 2017gfo, some SGRBs have been associated with kilonova candidates, such as GRB 050709 (Jin et al., 2016), GRB 130603B (Berger et al., 2013; Tanvir et al., 2013) and GRB 160821B (Lü et al., 2017; Lamb et al., 2019; Troja et al., 2019). However, from an observational point of view, it is not only SGRBs that exhibit the features of the merger, but also several LGRBs, which are more likely SGRBs with extended emission (EE-SGRB), showing potential kilonova signatures but lacking an associated supernova. For example, GRB 060614 (Della Valle et al., 2006; Fynbo et al., 2006; Gal-Yam et al., 2006; Gehrels et al., 2006; Metzger et al., 2008; Yang et al., 2015), GRB 211211A (Rastinejad et al., 2022; Troja et al., 2022; Yang et al., 2022; Chang et al., 2023; Gompertz et al., 2023), GRB 211227A (Lü et al., 2022; Ferro et al., 2023), and GRB 230307A (Gillanders et al., 2023; Du et al., 2024; Levan et al., 2024; Yang et al., 2024; Sun et al., 2025). On the contrary, the short GRB 200826A, which seems to be associated with an SN, suggests an origin from the collapse of massive stars (Ahumada et al., 2021; Zhang et al., 2021). In summary, the traditional long versus short GRB scheme no longer always corresponds to the two distinct physical origins (Zhang, 2006; Lü et al., 2010; Berger et al., 2013; Lü et al., 2014).
@@
One interesting question focuses on the EE-SGRBs, the BNS mergers and NS-BH mergers are significantly challenged in explaining the duration of their prompt $\gamma$-ray emission due to their shorter timescales of the mergers. Thus, the neutron star-white dwarf (NS-WD) merger has been proposed to explain the observational properties of EE-SGRBs (Yang et al., 2022; Zhong et al., 2023; Wang et al., 2024). The reason is that WDs are less dense than NSs, and the free-fall timescale of NS-WD mergers could be significantly longer than that of BNS mergers or NS-BH mergers. If this is the case, it could naturally explain the long timescale characteristics of these short GRBs with extended emission. However, WDs are proton-rich, so the ejecta from the NS-WD mergers may not be as neutron-rich as those from systems like NS-NS mergers, NS-BH mergers, Collapsar, etc. Therefore, the astrophysical environment of the NS-WD mergers seems not to be suitable for r-process elements production (Fernández et al., 2019; Bobrick et al., 2022; Kaltenborn et al., 2023).
@@
To understand the process of NS-WD mergers, we provide a relatively clear picture of NS-WD mergers: NS-WD mergers are originated from binaries formed through common envelope evolution and gravitational wave-driven orbital decay. Population synthesis suggests that such systems are abundant in the Galaxy (Nelemans et al., 2001). When unstable mass transfer begins, the WD is tidally disrupted to form an accretion disk around the NS (Papaloizou & Pringle, 1984; Fryer et al., 1999). Numerical simulations have extensively explored the subsequent disk evolution, which involves nuclear burning, fallback accretion, angular momentum transport, and outflows driven by neutrino cooling or magnetic processes. These features are somewhat similar to that of found in post-merger disks from NS-NS systems (Fernández & Metzger, 2013). Recently works have further investigated how the structure, composition, and thermodynamic state of the disk affect the ejecta mass, nucleosynthesis, and the properties of resulting faint and rapidly evolving transients (Zenati et al., 2019, 2020; Bobrick et al., 2022; Kaltenborn et al., 2023). Especially, Chen & Liang (2024) provides insight into the radioactive heating of heavy elements from nuclear burning, which may inform further studies of the NS-WD merger ejecta.
@@
In this work, we investigate the heavier elements produced by NS-WD mergers based on nucleosynthesis simulations. Then, we explore the luminosity of the corresponding kilonova-like emission produced by the radioactive decay of the elements produced by nucleosynthesis based on detailed numerical calculations. Finally, we compare the calculated luminosity of the kilonova-like emission with the observations of the possible kilonova associated with GRB 211211A. This paper is organized as follows. In Section 2, we introduce the details of nucleosynthesis and the method we adopted to calculate the kilonova-like light curves. The results of the nucleosynthesis elements and the corresponding kilonova-like light curves are presented in Section 3. The application to a possible kilonova emission associated with GRB 211211A based on our results is shown in Section 4. Conclusions are drawn in Section 5, along with some additional discussion. Throughout this paper, we adopt a concordance cosmology with the parameters $H_{\rm 0}=71~{}\rm km~{}s^{-1}~{}Mpc^{-1}$, $\Omega_{\rm M}=0.30$, and $\Omega_{\Lambda}=0.70$.
@@ 2 Methods
In general, based on the mass of white dwarfs ($M_{\rm WD}$) in WD-NS binaries, there are two main evolutionary pathways that are widely discussed for such binary systems, namely stable mass transfer ($M_{\rm WD}<M_{\rm WD,c}$) and unstable mass transfer ($M_{\rm WD}>M_{\rm WD,c}$), where $M_{\rm WD,c}\sim 0.2M_{\odot}$ is the critical mass of the WD (Margalit & Metzger, 2016). More than 20 WD-NS binaries have been discovered in our Galaxy, and the mass of WDs in such binaries is larger than the critical mass (O’Shaughnessy & Kim, 2010). It suggests that the majority of WD-NS binary systems will undergo a merger with the process of unstable mass transfer within a Hubble time (Toonen et al., 2018). In our work, we adopt the unstable mass transfer model, and select three different groups of WD-NS systems for simulation, with mass groups of (0.75-1.4) $M_{\odot}$, (1-1.25) $M_{\odot}$, and (1-1.8) $M_{\odot}$ which named as Model-A, Model-B, and Model-C, respectively (see Table 1).
@@ 2.1 Nucleosynthesis
The merger of binary compact stars (e.g., NS-NS, NS-BH) or supernova explosions is an important site for the nucleosynthesis of heavy elements. One interesting question is how heavy the heaviest element can be formed during the WD-NS mergers. In this Section, we perform the nucleosynthesis via the nuclear reaction network code SkyNet to do the numerical simulations (Lippuner & Roberts, 2015, 2017).
SkyNet can compute the evolution of nucleosynthesis in a wide range of astrophysical scenarios where nucleosynthesis occurs, and it has been widely used to evolve the abundance of nuclear species under the influence of nuclear reactions. The network contains 7843 nuclides and 140,000 reactions, ranging from free neutrons ($Z=0$) and protons ($Z=1$) to $\rm{{}^{337}Cn}$ ($Z=112$). SkyNet adopts a modified version of the Helmholtz equation of state (EOS; Timmes & Swesty (2000)) to calculate the entropy. The nuclear reaction rates used in SkyNet are taken from the JINA REACLIB database (Cyburt et al., 2010). Nuclear masses and partition functions are included in the WebNucleo XML file distributed with REACLIB for nuclide species that are without experimental data, and we use the theoretical data derived from the finite-range droplet model (FRDM; Möller et al. (2015)).
The first step is to evolve the accretion disk, which is composed of material generated by the tidal disruption of the WD. The initial density profile of the disk that we adopt can be described in Margalit & Metzger (2016), and we adopt the average density as the initial condition density in our simulations. In terms of the temperature, we investigate the effect of different initial temperatures due to the poorly known temperature. Thus, three groups of temperatures, such as 4 GK, 5 GK, and 6 GK, are used in the simulations. We initialize the composition of the material in the disk, namely, equal parts of carbon and oxygen, and with 1% helium. During the disk evolution, we assume that the overall density and temperature of the whole disk do not change too much, and we adopt a constant density and temperature trajectory in SkyNet to perform the simulations. The total timescale of disk evolution is set to 60 seconds, and one can obtain the final abundance of the disk evolution.
Then, we set the final abundance of the disk evolution as the initial abundance of the ejecta evolution. In our nucleosynthesis simulations of the ejecta, for the density profile we use the parameterized trajectory provided by Lippuner & Roberts (2015). The initial density of the ejecta is set to $\rho_{\rm ej}=3.00\times 10^{5}\rm~{}g/cm^{3}$ for Model-A, which is taken from the results of Zenati et al. (2019). By assuming that the ejecta of Model-B and Model-C share the same radius as in Model-A, together with the results of numerical relativity simulations given by Kaltenborn et al. (2023), one can derive the initial densities of Model-B and Model-C to be $4.42\times 10^{5}\rm g/cm^{3}$ and $3.68\times 10^{5}\rm g/cm^{3}$, respectively. Table 1 shows the details of the initial parameter settings for both disk and ejecta. Moreover, the expansion timescale is approximately set to be 3 seconds, considering the long duration of WD-NS mergers. For ejecta evolution, we also set three groups of initial temperatures, which are 4 GK, 5 GK, and 6 GK, respectively. Then, the temperature profiles are set as $TV^{1/3}=\rm const$ for the evolution of the ejecta, which is because the expansion of the ejecta is nearly adiabatic (Wollaeger et al., 2018; Farouqi et al., 2010).
@@ 2.3 Kilonova-like emission
The kilonova emission, which is an optical/infrared transient, is generated in the ejected material and is powered by the radioactive decay of r-process nuclei (Li & Paczyński, 1998; Metzger et al., 2010; Yang et al., 2015; Ma et al., 2018). If extra energy is injected into such a transient, excepting radioactive decay of r-process nuclei, the kilonova will be brighter, called merger-nova (Yu et al., 2013; Gao et al., 2017; Yuan et al., 2021). In our work, the merger ejecta is assumed to be spherically symmetric (Kasen et al., 2017), even some numerical simulations of double NS merger suggest that the ejecta or radioactive decay of r-process are more or less anisotropic (Ciolfi et al., 2017; Siegel & Metzger, 2018; Collins et al., 2023). By adopting an average velocity of ejecta $v_{\rm ej}=0.1$ c taken from the simulation result of Zenati et al. (2019), and following the method in Chen & Liang (2024), we divide the merger ejecta into $n$ layers, and treat the evolution of each layer independently. The velocities of the innermost layer (first layer) and the outermost layer (last layer) are set to $v_{\rm min}=v_{\rm ej}-0.03c$ and $v_{\rm max}=v_{\rm ej}+0.03c$, respectively. Thus, the expansion velocity of the $i$th layer is given by
@@
where $E_{i}$ is the internal energy of the $i$th layer, $\dot{Q}(t)$ is the heating rate, which can be given by Eq.(1). $L_{i}$ is the radiation luminosity of each layer, which can be estimated as
@@
where $t_{d,i}=\tau_{i}v_{i}t/c$ is the photon diffusion timescale, and $t_{lc,i}=v_{i}t/c$ is the energy loss time limitation to the light crossing time. The optical depth ($\tau_{i}$) of the $i$th shell is given by
@@
where $\sigma_{\rm SB}$ is the Stefan-Boltzmann constant, $R_{\rm ph}$ is the photospheric radius when the optical depth equals unity, i.e.,

### v2

@@ Abstract
Long-duration GRB 211211A, which lacked an associated supernova at very a low redshift $z=0.076$ but was associated with a possible kilonova emission, has attracted great attention. The neutron star-white dwarf (NS-WD) merger is proposed as a possible progenitor of GRB 211211A, and it could naturally explain the long duration of the prompt emission. However, the NS-WD merger is not an ideal site for producing heavy elements via r-process nucleosynthesis. In this Letter, we investigate the heavy elements produced in NS-WD mergers based on numerical simulations of nucleosynthesis via SkyNet and then calculate the resulting kilonova-like emission to compare with the solidly observed case of possible kilonova emission associated with GRB 211211A. By adopting three models (i.e., Model-A, Model-B, and Model-C) from M. A. R. Kaltenborn et al. at different temperatures ($T=4$ GK, 5 GK, and 6 GK), which are treated as free parameters, we find that the mass number of the heaviest element produced in our simulations is less than 90 ($A<90$). Moreover, by comparing the calculated kilonova-like emission with the afterglow-subtracted observations of the possible kilonova associated with GRB 211211A, it is found that the merger of an NS and WD cannot be ruled out as the origin of GRB 211211A to produce the possible kilonova emission if the remnant of the WD-NS merger is a supramassive or stable magnetar. Otherwise, it is difficult to explain the early possible kilonova emission following GRB 211211A by invoking the merger of a WD and an NS.
Gamma-ray bursts
@@ 1 Introduction
Gamma-ray bursts (GRBs), the most luminous electromagnetic explosions in the Universe, are traditionally classified into two categories based on their duration: short-duration GRBs (SGRBs; $T_{90}<2s$) and long-duration GRBs (LGRBs; $T_{90}>2s$) (Kouveliotou et al., 1993). Robust associations of the underlying supernovae (SNe) with some LGRBs suggest that they are likely related to the collapse of massive stars (Woosley, 1993; Galama et al., 1998; Stanek et al., 2003; Fruchter et al., 2006; Pian et al., 2006; Woosley & Bloom, 2006). In contrast, some SGRBs are associated with nearby host galaxies with little star formation, and others do not have an associated SN (Berger et al., 2005; Fox et al., 2005; Hjorth et al., 2005; Tanvir et al., 2005), suggesting that the SGRBs are possibly originated from the merger of binary compact stars, such as binary neutron star (BNS) mergers (Paczynski, 1986; Eichler et al., 1989) and neutron star-black hole (NS-BH) mergers (Paczynski, 1991). On the other hand, such binary systems are also considered to be potential sources of gravitational-wave radiation and kilonova emission (Berger, 2014; Zhang, 2018). The first “smoking gun” evidence to confirm the BNS merger as the source of SGRBs, gravitational-wave radiation, and kilonova, emerged from the direct detection of GW 170817 via Advanced LIGO and Virgo, which was associated with electromagnetic counterparts, such as GRB 170817A and AT 2017gfo (Abbott et al., 2017; Alexander et al., 2017; Arcavi et al., 2017; Chornock et al., 2017; Coulter et al., 2017; Covino et al., 2017; Drout et al., 2017; Evans et al., 2017; Goldstein et al., 2017; Kasen et al., 2017; Kasliwal et al., 2017; Pian et al., 2017; Savchenko et al., 2017; Shappee et al., 2017; Smartt et al., 2017; Tanaka et al., 2017; Tanvir et al., 2017; Zhang et al., 2018).
@@
Except for the confirmed event GW 170817/ GRB 170817A/ AT 2017gfo, some SGRBs have been associated with kilonova candidates, such as GRB 050709 (Jin et al., 2016), GRB 130603B (Berger et al., 2013; Tanvir et al., 2013) and GRB 160821B (Lü et al., 2017; Lamb et al., 2019; Troja et al., 2019). However, from an observational point of view, it is not only SGRBs that exhibit the features of the merger but also several LGRBs, which are more likely SGRBs with extended emission (EE-SGRB), showing potential kilonova signatures but lacking an associated SN, for example, GRB 060614 (Della Valle et al., 2006; Fynbo et al., 2006; Gal-Yam et al., 2006; Gehrels et al., 2006; Metzger et al., 2008; Yang et al., 2015), GRB 211211A (Rastinejad et al., 2022; Troja et al., 2022; Yang et al., 2022; Chang et al., 2023; Gompertz et al., 2023), GRB 211227A (Lü et al., 2022; Ferro et al., 2023), and GRB 230307A (Gillanders et al., 2023; Du et al., 2024; Levan et al., 2024; Yang et al., 2024; Sun et al., 2025). On the other hand, the short GRB 200826A, which seems to be associated with an SN, suggests an origin from the collapse of massive stars (Ahumada et al., 2021; Zhang et al., 2021). In summary, the traditional long-versus-short-GRB scheme no longer always corresponds to the two distinct physical origins (Zhang, 2006; Lü et al., 2010; Berger et al., 2013; Lü et al., 2014).
@@
One interesting question focuses on the EE-SGRBs; it is challenging for the BNS mergers and NS-BH mergers to explain the duration of their prompt $\gamma$-ray emission due to their shorter timescales of the mergers. Thus, the neutron star-white dwarf (NS-WD) merger has been proposed to explain the observational properties of EE-SGRBs (Yang et al., 2022; Zhong et al., 2023; Wang et al., 2024). The reason is that WDs are less dense than NSs, and the free-fall timescale of NS-WD mergers could be significantly longer than that of BNS mergers or NS-BH mergers. If this is the case, it could naturally explain the long timescale characteristics of these EE-SGRBs. However, WDs are proton rich, so the ejecta from the NS-WD mergers may not be as neutron rich as those from systems like NS-NS mergers, NS-BH mergers, collapsars, etc. Therefore, the astrophysical environment of the NS-WD mergers seems not to be suitable for r-process element production (Fernández et al., 2019; Bobrick et al., 2022; Kaltenborn et al., 2023).
@@
To understand the process of NS-WD mergers, we provide a relatively clear picture of NS-WD mergers: NS-WD mergers are originated from binaries formed through common envelope evolution and gravitational-wave-driven orbital decay. Population synthesis suggests that such systems are abundant in the Galaxy (Nelemans et al., 2001). When unstable mass transfer begins, the WD is tidally disrupted to form an accretion disk around the NS (Papaloizou & Pringle, 1984; Fryer et al., 1999). Numerical simulations have extensively explored the subsequent disk evolution, which involves nuclear burning, fallback accretion, angular momentum transport, and outflows driven by neutrino cooling or magnetic processes. These features are somewhat similar to that of found in postmerger disks from NS-NS systems (Fernández & Metzger, 2013). Recent works have further investigated how the structure, composition, and thermodynamic state of the disk affect the ejecta mass, nucleosynthesis, and the properties of the resulting faint and rapidly evolving transients (Zenati et al., 2019, 2020; Bobrick et al., 2022; Kaltenborn et al., 2023). Especially, Chen & Liang (2024) provide insight into the radioactive heating of heavy elements from nuclear burning, which may inform further studies of the NS-WD merger ejecta.
@@
In this work, we investigate the heavier elements produced by NS-WD mergers based on nucleosynthesis simulations. Then, we explore the luminosity of the corresponding kilonova-like emission produced by the radioactive decay of the elements produced by nucleosynthesis based on detailed numerical calculations. Finally, we compare the calculated luminosity of the kilonova-like emission with the observations of the possible kilonova associated with GRB 211211A. This Letter is organized as follows. In Section 2, we introduce the details of nucleosynthesis and the method we adopted to calculate the kilonova-like light curves. The results of the nucleosynthesis elements and the corresponding kilonova-like light curves are presented in Section 3. The application to a possible kilonova emission associated with GRB 211211A based on our results is shown in Section 4. Conclusions are drawn in Section 5, along with some additional discussion. Throughout this Letter, we adopt a concordance cosmology with the parameters $H_{\rm 0}=71~\rm km~s^{-1}~Mpc^{-1}$, $\Omega_{\rm M}=0.30$, and $\Omega_{\Lambda}=0.70$.
@@ 2 Methods
In general, based on the mass of white dwarfs ($M_{\rm WD}$) in WD-NS binaries, there are two main evolutionary pathways that are widely discussed for such binary systems, namely stable mass transfer ($M_{\rm WD}<M_{\rm WD,c}$) and unstable mass transfer ($M_{\rm WD}>M_{\rm WD,c}$), where $M_{\rm WD,c}\sim 0.2M_{\odot}$ is the critical mass of the WD (Margalit & Metzger, 2016). More than 20 WD-NS binaries have been discovered in our Galaxy, and the mass of WDs in such binaries is larger than the critical mass (O’Shaughnessy & Kim, 2010). It suggests that the majority of WD-NS binary systems will undergo a merger with the process of unstable mass transfer within a Hubble time (Toonen et al., 2018). In our work, we adopt the unstable mass transfer model and select three different groups of WD-NS systems for simulation, with mass groups of (0.75-1.4) $M_{\odot}$, (1-1.25) $M_{\odot}$, and (1-1.8) $M_{\odot}$, which we named as Model-A, Model-B, and Model-C, respectively (see Table 1).
@@ 2.1 Nucleosynthesis
The merger of binary compact stars (e.g., NS-NS, NS-BH) or SN explosions is an important site for the nucleosynthesis of heavy elements. One interesting question is how heavy the heaviest element can be formed during the WD-NS mergers. In this section, we perform the nucleosynthesis via the nuclear reaction network code SkyNet to do the numerical simulations (Lippuner & Roberts, 2015, 2017).
SkyNet can compute the evolution of nucleosynthesis in a wide range of astrophysical scenarios where nucleosynthesis occurs, and it has been widely used to evolve the abundance of nuclear species under the influence of nuclear reactions. The network contains 7843 nuclides and 140,000 reactions, ranging from free neutrons ($Z=0$) and protons ($Z=1$) to $\rm{{}^{337}Cn}$ ($Z=112$). SkyNet adopts a modified version of the Helmholtz equation of state (EOS; Timmes & Swesty (2000)) to calculate the entropy. The nuclear reaction rates used in SkyNet are taken from the JINA REACLIB database (Cyburt et al., 2010). Nuclear masses and partition functions are included in the WebNucleo XML file distributed with REACLIB for nuclide species that are without experimental data, and we use the theoretical data derived from the finite-range droplet model (Möller et al., 2015).
The first step is to evolve the accretion disk, which is composed of material generated by the tidal disruption of the WD. The initial density profile of the disk that we adopt can be described in Margalit & Metzger (2016), and we adopt the average density as the initial condition density in our simulations. In terms of the temperature, we investigate the effect of different initial temperatures due to the poorly known temperature. Thus, three groups of temperatures, such as 4, 5, and 6 GK, are used in the simulations. We initialize the composition of the material in the disk, namely, equal parts of carbon and oxygen, and with 1% helium. During the disk evolution, we assume that the overall density and temperature of the whole disk do not change too much, and we adopt a constant density and temperature trajectory in SkyNet to perform the simulations. The total timescale of disk evolution is set to 60 s, and one can obtain the final abundance of the disk evolution.
[…]
@@ 2.3 Kilonova-like emission
The kilonova emission, which is an optical/infrared transient, is generated in the ejected material and is powered by the radioactive decay of r-process nuclei (Li & Paczyński, 1998; Metzger et al., 2010; Yang et al., 2015; Ma et al., 2018). If extra energy is injected into such a transient, except for radioactive decay of r-process nuclei, the kilonova will be brighter and is called merger-nova (Yu et al., 2013; Gao et al., 2017; Yuan et al., 2021). In our work, the merger ejecta is assumed to be spherically symmetric (Kasen et al., 2017); even some numerical simulations of double NS merger suggest that the ejecta or radioactive decay of r-process is more or less anisotropic (Ciolfi et al., 2017; Siegel & Metzger, 2018; Collins et al., 2023). By adopting an average velocity of ejecta $v_{\rm ej}=0.1$ c taken from the simulation result of Zenati et al. (2019), and following the method in Chen & Liang (2024), we divide the merger ejecta into $n$ layers and treat the evolution of each layer independently. The velocities of the innermost layer (first layer) and the outermost layer (last layer) are set to $v_{\rm min}=v_{\rm ej}-0.03c$ and $v_{\rm max}=v_{\rm ej}+0.03c$, respectively. Thus, the expansion velocity of the $i$th layer is given by
@@
where $E_{i}$ is the internal energy of the $i$th layer and $\dot{Q}(t)$ is the heating rate, which can be given by Equation (1). $L_{i}$ is the radiation luminosity of each layer, which can be estimated as
@@
where $t_{d,i}=\tau_{i}v_{i}t/c$ is the photon diffusion timescale and $t_{lc,i}=v_{i}t/c$ is the energy loss time limitation to the light-crossing time. The optical depth ($\tau_{i}$) of the $i$th shell is given by
@@
where $\sigma_{\rm SB}$ is the Stefan-Boltzmann constant and $R_{\rm ph}$ is the photospheric radius when the optical depth equals unity, i.e.,